# Evaluation of LLMs for mathematical problem solving

### v1

@@ Abstract
Large Language Models (LLMs) have shown impressive performance on a range of educational tasks, but are still understudied for their potential to solve mathematical problems. In this study, we compare three prominent LLMs, including GPT-4o, DeepSeek-V3, and Gemini-2.0, on three mathematics datasets of varying complexities (GSM8K, MATH500, and UNSW datasets). We take a five-dimensional approach based on the Structured Chain-of-Thought (SCoT) framework to assess final answer correctness, step completeness, step validity, intermediate calculation accuracy, and problem comprehension. The results show that GPT-4o is the most stable and consistent in performance across all the datasets, but particularly it performs outstandingly in high-level questions of the UNSW dataset. DeepSeek-V3 is competitively strong in well-structured domains such as optimisation, but suffers from fluctuations in accuracy in statistical inference tasks. Gemini-2.0 shows strong linguistic understanding and clarity in well-structured problems but performs poorly in multi-step reasoning and symbolic logic. Our error analysis reveals particular deficits in each model: GPT-4o is at times lacking in sufficient explanation or precision; DeepSeek-V3 leaves out intermediate steps; and Gemini-2.0 is less flexible in mathematical reasoning in higher dimensions.
@@ 3 Methodology
We first present the framework used to evaluate selected LLMs for mathematical problem-solving. Our approach combines multiple benchmark datasets along with UNSW School of Mathematics and Statistics assignments and exams. We incorporate controlled prompting strategies, with both quantitative and qualitative evaluation.
@@ 3.1 Mathematical problems
To ensure representativeness and reliability of the evaluation results, this study selected three sets of mathematical problems from different levels and backgrounds: GSM8K, MATH 500, and the graduate course question bank of the University of New South Wales (UNSW). These sets cover a wide range of problem types, from basic arithmetic and algebraic operations to advanced statistical reasoning and optimisation, and are characterised by diverse structures and distinct language styles.
@@
    UNSW Graduate Question Bank: This dataset consists of exam and assignment questions from the UNSW School of Mathematics and Statistics, covering probability inference, matrix decomposition, optimisation theory, and other courses. The language is rigorous, and the expression style is academic. Some questions have open structures, conceptual explanations, and non-standard expressions. The details about this dataset can be shown in Table 2.
@@
UNSW Problem 130 Advanced UNSW course materials
@@
Table 2: Summary of mathematical problem dataset from UNSW School of Mathematics and Statistics.
Course Code & Title Assessment Type Year Difficulty Level
MATH5335 Mathematical Computing Methods for Finance Final Exam 2013 Advanced/Postgraduate
Final Exam 2015 Advanced/Postgraduate
MATH5165 Optimisation Class Test 1 2024 Advanced/Undergraduate/Postgraduate
Class Test 2 2024 Advanced/Undergraduate/Postgraduate
MATH5905 Statistical Inference Assignment 1 2024 Postgraduate
Assignment 2 2024 Postgraduate
@@ 3.2 Automated LLM Evaluation Framework
    UNSW Problem dataset (real academic assessment scenarios (Table 2))
@@
    Stage 2 - Dataset pre-processing: We select many mathematical questions from the GSM8K, MATH500, and UNSW datasets (Table 1), covering multiple levels from basic arithmetic to advanced reasoning. After pre-processing, we classify the questions according to knowledge points and difficulty levels to facilitate subsequent layered evaluation and error analysis. We need to balance question types to ensure that different kinds of problems are represented in the evaluation.
@@
    UNSW A quantitative analysis was conducted on 130 mathematical problems from various subjects at the University of New South Wales to study the performance of language models in advanced mathematics.
@@ 3.3 Evaluation Metrics and LLM Prompting Strategies
The LLM evaluation focuses on two aspects: algebraic expression computation and mathematical reasoning problems [119]. Since LLMs are generative models, ensuring that the generated answers are accurate becomes very tricky, especially for mathematical reasoning tasks [120]. In the case of computational problems, we prefer to provide standard reference answers so that we can easily compare the model-generated answers to the correct answers. This allows us to evaluate the model’s performance with numerical precision [121]. We adopt the automatic proof verification method for the mathematical proof problem. Instead of relying on manual evaluations, we test the performance of the model by establishing the SCoT framework and an automatic scoring method inspired by recent advances in LLM-based reasoning verification [ren2023self]. It provides a more objective criterion for evaluating mathematical reasoning problems and improves the efficiency of the evaluation process [122]. We also provide a prompting mechanism to assess the ability to correct their mistakes. If a model provided an incorrect answer, it would be prompted again to solve the problem, and the corrected answer would be evaluated.
@@ 4.1 LLMs for solving GSM8K problems
We investigated more about the errors contributing to this mediocre result, including interruptions in the reasoning chain, omissions of steps, misunderstandings of quantitative relationships, and errors in unit conversion. In the case of GSM8K, all three models sometimes arrives at the correct answer. However, they cannot generate intermediate steps due to an unstable reasoning process or improper handling of details, which reflects their lack of coherence and robustness in multi-step tasks, as shown in Figure 8.
@@
Table 5 shows that Gemini performs relatively best on the GSM8K dataset, with an accuracy of 53.9%. In addition, it demonstrates a more complete reasoning path compared to GPT-4o (50.0%) and DeepSeek-V3 (42.6%)  [^2] https://github.com/SHARP-Mitsuko/Evaluation-of-LLMs-for-mathematical-problem-solving/tree/main/result/DeepSeek. To better understand this difference, we qualitatively (manually) review representative model responses, as summarised in Table 6. Gemini produced fewer instances of step skipping, unit confusion, and misinterpretation of the question intent. In contrast, we also found that DeepSeek’s reasoning chain is the most unstable, with frequent errors in unit conversion and logical expression, and is at the bottom of the three in terms of stability and accuracy.
@@ 4.2 LLMs for MATH500
4.3 LLMs for UNSW dataset
Table 10: Accuracy (%) of LLMs for UNSW Datasets
Model GPT-4o DeepSeek Gemini
Accuracy 35.6% 42.8% 54.4%
We evaluated LLM performance on the UNSW mathematics dataset, building on the qualitative analysis of incorrect answers using a structured scoring rubric as shown in Figure 1. We designed this rubric to assess the quality of mathematical reasoning, presentation, and accuracy with criteria shown in Table 11:
Table 11: The Score Criteria
Score Criteria
1 A completely incorrect solution with major logical flaws or missing fundamental steps
2 A weak understanding with multiple errors or incoherent reasoning
3 The main method is correct, but contains explanation or calculation problems
4 A mostly correct solution with minor mistakes or slightly informal reasoning
5 Fully correct, well-organised, and rigorous answers that represent a model solution
We excluded all sub-questions that received a full score of 5. For each non-perfect response, we conducted detailed error analysis to identify the primary cause of the mistake. We assigned each response to only one dominant error category to ensure consistency, even if multiple issues were present. This approach formed the foundation for our subsequent analysis of error types. We developed a taxonomy of ten categories to describe and classify the nature of the errors, which reflect different aspects of failure in LLM-generated reasoning as shown in Table 14.
Table 12: LLM performance based on scores on UNSW problems by subject, where higher scores indicate higher correctness of solutions.
Subject Score DeepSeek Gemini GPT-4o
Optimisation 1 6.7% 0.0% 0.0%
2 6.7% 0.0% 0.0%
3 20.0% 0.0% 0.0%
4 6.7% 20.0% 53.0%
5 60% 80% 46.7%
Computational Methods for Finance 1 19.0% 2.6% 0.0%
2 10.1% 6.5% 1.2%
3 10.1% 31.2% 8.8%
4 10.1% 40.3% 42.5%
5 50.6% 19.5% 47.5%
Statistical Inference 1 33.3% 12.8% 0.0%
2 35.9% 25.8% 0.0%
3 7.7% 17.9% 7.7%
4 5.1% 15.4% 33.3%
5 17.9% 28.2% 59.0%
Table 12 shows that the GPT-4o score distribution differs significantly between subjects in the UNSW dataset. In Statistical Inference, the model achieves the highest proportion of perfect scores (59.0%). This indicates strong performance in problems involving standard statistical procedures such as MLE (Maximum Likelihood Estimation). This suggests that GPT-4o is well-suited to problems with familiar mathematical structures and clear solution paths. In optimisation, GPT-4o has the highest proportion of 4-point responses (53.3%), with slightly fewer perfect scores (46.7%). This may reflect that while GPT-4o can follow correct procedures, such as computing gradients, it often lacks full formal justification and presentation, which prevents it from achieving a perfect score under strict grading criteria. In Computational Methods for Finance, the model has the highest rate of low scores (3 and below, totalling around 10%). These errors may stem from difficulties interpreting domain-specific terminology, probabilistic reasoning, or less familiar symbolic conventions. Overall, GPT-4o’s performance depends heavily on problem type, with stronger results on structurally straightforward and procedurally familiar tasks.
Table 13: Mainly Shortcoming Details
Type Description
Understanding Errors Misunderstanding of key concepts, such as confusion between prior, posterior, and MLE
Deduction Errors Correct understanding of concepts, but errors in formula derivation or transformation steps
Expression Errors Correct calculations or reasoning, but unclear writing, lack of explanation, or confusing symbols
Process Control Errors Disorganised problem-solving process, inconsistent answers to sub-questions, or missing conditions
Format Errors Minor errors that do not affect the core logic, such as copying formulas incorrectly, writing variables incorrectly, or missing symbols
The SCoT framework facilitates a clearer evaluation of LLM performance by decomposing their reasoning into structured steps. The average score of GPT-4o performs well in MATH5905 Statistical Inference course (see Table 16) mainly because it follows standard statistical procedures and generates coherent solutions that align with the SCoT criteria, particularly in terms of logical completeness and consistency. This matches the high proportion of perfect scores in MATH5905 Statistical Inference, showing that GPT-4o performs well when the problem is clear and the reasoning is well-defined. We designed a prompt that guides GPT-4o through solving each sub-question step-by-step, and an automated grading model based on Deepseek-V3 that evaluates each solution step. The LLM returns a structured JSON format answer for each step which allows us to check the logic and correctness at every stage, check the model’s reasoning, and easily spot gaps. This step-by-step evaluation ensures that we can detect issues in the model’s reasoning, such as missing intermediate steps, inaccurate calculations, or insufficient explanations.Table 14 shows the types of errors made by GPT-4o on the UNSW dataset. In addition, we also reviewed and categorised the errors manually from the model’s results that did not receive full marks, helping to highlight common mistakes in multi-step mathematical problem solving with the following:
  • 1.
    Lack of Justification or Explanation (39%)
  • 2.
    Less Precise (33%)
  • 3.
    Missing Intermediate Steps or Details (9%)
Table 14: Main shortcomings of answers on UNSW dataset
Error type Model Percent
Lack of justification or explanation GPT 39%
DeepSeek 21%
Gemini 20%
Less precise GPT 33%
DeepSeek 3%
Gemini 3%
Missing intermediate steps or details GPT 9%
DeepSeek 24%
Gemini 18%
Unclear or in consistent notation GPT 8%
DeepSeek 7%
Gemini 18%
Poor structure or lack of formal presentation GPT 3%
DeepSeek 14%
Gemini 17%
Incomplete gradient/Hessian analysis GPT 3%
DeepSeek 3%
Gemini 3%
Lack of logical flow GPT 3%
DeepSeek 9%
Gemini 2%
Missing eigenvalue/matrix-related reasoning GPT 2%
DeepSeek -
Gemini 7%
Verbose GPT -
DeepSeek 6%
Gemini 9%
These three types of errors account for over 80% of all observed errors(see Table 14), which indicates that GPT-4o can often identify the correct approach or final answer, but it frequently omits essential reasoning or fails to maintain computational precision. At the same time, the remaining error types related to notation clarity, logical coherence, and incomplete second-order analysis in optimisation problems which were observed less frequent but still noteworthy, This findings are based on manual review of model responses in the UNSW dataset, as summarised in Table 14.
In general, most of GPT-4o’s errors (Table 14) are mainly due to insufficient formal expressions rather than a lack of understanding of mathematical concepts. Therefore, introducing stronger logical constraints and formal specifications in the prompt design is expected to substantially improve the rigour of model reasoning and problem-solving completeness. These findings suggest that improving the model’s ability to intermediate reasoning and clarity of its logical progression may significantly improve GPT-4o’s overall performance on mathematical tasks.
In the case of DeekSeek (Table 12), the distribution of scores is quite extreme, mainly distributed in Score 1, Score 2, and Score 5, with fewer middle scores. This indicates that DeepSeek either answers the questions perfectly or has many errors, and there are fewer incomplete answers. The reason for this phenomenon may be that DeepSeek specialises in specific fields, or it may be due to the underlying logic of the code itself that led to this result. Furthermore, we can observe that DeepSeek performs much better in computational methods and optimisation than in statistical inference. This also indicates that DeepSeek has an advantage in handling computational and optimisation problems, but is slightly inferior in analysis, statistics, and explanation, and can only solve relatively simple problems. Furthermore, by analysing the main shortcomings of answers, as shown in Table 14, it can be found that the most common shortcomings that DeepSeek encounters when answering are missing intermediate steps or details and a lack of justification or explanation. This is mainly manifested in DeepSeek providing answers with only partial methods, explanations, and conclusions, but lacking intermediate processes and detailed computational details. However, this may be related to the difficulty of the question, as relatively simple questions usually have a more detailed process, while more difficult questions may omit a large number of steps.
The distribution of scores shows that Gemini performs well in multi-step calculations and flexible application ofspecificn formulas, Table 12. The scores of answers in solving problems related to statistical inference are very close, which indicates that a mix of formulas and analysis can be a challenge for Gemini. Issues like these require solvers to tackle subsequent questions based on the results you derived from the former ones. Even if Gemini can retain all the context, it still cannot realise and correct previous mistakes in solving the rest of the problems. The problem-solving process lacks self-censorship; hence, the more subproblems there are, the higher the probability of getting a low score.
The shortcomings of Gemini-2.0, as shown in Table 14 highlight verbosity as a significant drawback where Gemini’s answers often take a longer path to the destination compared with the correct solution. Strangely, verbosity and lack of explicit explanation of key theory or variables can appear in an answer at the same time. Students know what they have learned in this course and try to use the knowledge to complete the exam. However, Gemini knows almost everything about this topic from a wide range of sources and it cannot figure out which key formula or concept is being examined in this problem  [^3] https://github.com/SHARP-Mitsuko/Evaluation-of-LLMs-for-mathematical-problem-solving. Gemini omits some main steps in the process of proof for the same reason. It takes some computational steps and interpretation for granted. However, such omissions can lead to a loss of marks when the answer is marked.
As an LLM, Gemini prefers to use natural language instead of mathematical language to explain conditions or variables. This is also why the answer is verbose, although using plain language is helpful for beginners to understand the core idea of this problem. The marker may get confused about the solutions expressed and lower the grade due to weak presentation, since Gemini chooses the notation freely and arbitrarily. Even though notation can vary across texts in different reference books, there is still a consensus on the usage of notation among students and scholars. Non-standard and uncommon notation may mislead the solver to a wrong result. The formatting issues mentioned above are not important if the final result is correct; however, it will make the answers ambiguous and less precise.
As shown in Table 14, the problems requiring proof and calculations are mainly included in the dataset across three fields, Optimisation, Computational Methods for Finance and Statistical Inference. Subjects selected are close to industrial practices, since the performance of LLMs has to be validated through real-world applications rather than pure mathematical theory. 29% accuracy is not much surprising due to the rigorous criterion. Only a full mark answer can be considered correct for proof problems.
4.4 Qualitative assessment for UNSW problems
We studied MATH5905 Statistical Inference LLM solutions from the UNSW dataset qualitatively to evaluate their capability in dealing with complex mathematical inference questions. The question consists of four sub-problems, which examine the determination of sufficiency and completeness, the unbiased verification of maximum likelihood estimation(MLE), the derivation of asymptotic distributions, and the application of multivariate delta methods, and involves several core concepts in statistical theory, with a rigorous logical structure and a coherent step-by-step approach, which makes it an ideal sample for evaluating the performance of LLMs to perform mathematical reasoning  [^4] https://github.com/SHARP-Mitsuko/Evaluation-of-LLMs-for-mathematical-problem-solving/blob/main/datasets/UNSW-problems/Stat2.pdf. The results of the study, as shown in the Table 15, show that the mainstream models usually identify basic solution strategies and can initially construct a reasoning framework. However, the models generally have problems in the specific derivation process. We take the completeness proof to provide an example. Although the model can identify sufficient statistics, it often fails to apply theoretical tools such as the Lehmann-Scheffé theorem; in the unbiasedness test, although part of the model can correctly judge the direction, it lacks the detailed development of the expectation calculation; in the application of the multivariate delta method, even though it is close to being correct in form, it lacks the matrix and parameter independence premise. Although the form is nearly correct in applying the multivariate Delta method, there is a lack of a matrix inversion step and parameter independence. These problems can be categorised as ’reasoning chain breaks’ or ’logical jumps’, especially in the case of a longer structure and higher logical relevance of sub-problems. The lack of rigour in the early reasoning of the model will be cumulatively amplified to the subsequent sub-problems, resulting in the so-called ’cumulative collapse’. The so-called ’cumulative breakdown’ occurs. In addition, dimensional analyses of the SCoT framework show that its output struggles to meet the basic requirements of logical closure and mathematical rigour for higher-order reasoning problems. These errors are also repeated to varying degrees in other UNSW questions, reflecting that there are still some stability and controllability challenges in formal mathematical tasks in the current LLM  [^5] https://github.com/SHARP-Mitsuko/Evaluation-of-LLMs-for-mathematical-problem-solving/tree/main/result.
Table 15: Summary of LLM’s performance on a four-part statistical inference problem
Sub-question Content LLMs Output Summary Common Issue
Q1 Sufficiency and completeness justification Identified exponential family structure correctly, but failed to invoke formal completeness arguments such as Lehmann–Scheffé theorem. Missing statistical justification
Q2 Unbiasedness of MLE for $1/(1+\theta)$ Provided correct final answer but omitted expected value calculation and skipped integration steps. Incomplete derivation
Q3 Asymptotic distribution (Delta Method) Applied Delta Method without deriving the transformation derivative or specifying Fisher Information. Lack of precision and justification
Q4 Multivariate Delta Method application General form used correctly, but failed to mention parameter independence or invert the Fisher Information matrix. Incomplete assumptions and structure
Note: The above summary is based on current models’ outputs. Further analysis will incorporate additional model results for a more comprehensive comparison.
4.5 Summary of results: LLMs across the three mathematical problem datasets
We finally present a summary of results to review the strengths and weaknesses of the respective LLMs across the three mathematical problem datasets, including GSM8K, MATH500 and UNSW datasets.
As shown in Table 17, the results show that the accuracy of GPT-4o, DeepSeek, and Gemini for the GSM8K dataset is 54.7%, 54.4%, and 53.9% , respectively. This is generally maintained at a medium-low level from 50% to 55%, and the difference between the models is small, among which GPT-4o has a slight advantage. GSM8K contains problems that require the LLMs to find the results step by step, which requires logical reasoning and simple calculation. GSM8K problems can be solved easily by college students and even high school students [113]. However, we find that the performance of LLMs on GSM8K tasks is average, and it seems that there is a 50 % chance of getting it right. The similar accuracy indicates that the problem does not lie in the difference between models, but in the limitation of logical reasoning within LLMs themselves, which makes them not suitable for this set of problems. Therefore, all three models can be helpful when facing problems like GSM8K, but they cannot guarantee correct answers.
The overall accuracy for UNSW problems is 29%, which is satisfactory in Table 17. The main reason for this accuracy is that a majority of answers receive scores of 3 or 4, especially problems from Computational Methods for Finance. DeepSeek has a much higher accuracy rate in handling UNSW problems (unlike MATH500), with an overall accuracy rate of 42.8%, Table 17.
Finally, we further summarise the overall performance of LLMs under different subjects in the UNSW dataset by computing the weighted average score of each LLM under each subject based on the distribution of scores in Table 12. We compute the average score $S$ by:
$S=\sum_{i=1}^{5}i\cdot p_{i}$ (4.1)
where $p_{i}$ denotes the proportion of the model receiving score i in that subject. As shown in Table 16, GPT-4o has a stable average score in several domains, especially in UNSW MATH5905 Statistical Inference, where it performs the best, with a score of 4.51. This reflecting its good adaptability to standard statistical tasks. Gemini shows a clear advantage in the MATH5165 Optimisation problem, with an average score of 4.80, but it still suffers from some redundancy in linguistic expression and notation specification. In contrast, the performance of DeepSeek-V3 varies greatly across subjects, with a fair score of 3.63 in Computational Methods for Finance, but a significantly weaker score of 2.38 in MATH5905 Statistical Inference, mainly due to an incomplete inference chain and missing intermediate steps. These weighted average scores provide quantitative support for differences in the capabilities of LLMs in domain-specific tasks and further corroborate what the previous section revealed based on the quality of the reasoning chain.
Table 16: Average Score of LLMs for UNSW Dataset by Subject
Course Code Subject DeepSeek Gemini GPT-4o
MATH5165 Optimisation 4.07 4.80 4.46
MATH5335 Computational Methods for Finance 3.63 3.68 4.36
MATH5905 Statistical Inference 2.38 3.21 4.51
Table 17: Overall Accuracy (%) of LLMs on Three Different Datasets
Model Dataset Accuracy
DeepSeek GSM8K 42.6%
MATH500 51.1%
UNSW 42.8%
Gemini-2.0 GSM8K 53.9%
MATH500 54.7%
UNSW 54.4%
GPT-4o GSM8K 50.0%
MATH500 36.8%
UNSW 35.6%
5 Discussion
We compared three LLMs, including GPT-4o, Gemini-2.0, and DeepSeek-V3—on mathematical problem-solving across varying difficulty levels. In addition to benchmark datasets, we used the UNSW dataset to assess models that feature diverse problem formats and non-standardised expressions, requiring both procedural precision and semantic comprehension. We observed that the LLMs can stably output complete answers containing intermediate variables and formula calculation steps when faced with moderately complex algebra, geometry, and probability questions. Furthermore, we wanted to evaluate whether the LLMs can automatically decompose sub-tasks and determine solution strategies, with relatively clear reasoning logic. Gemini-2.0 exhibited the most stable and structured reasoning patterns, particularly excelling in multi-step tasks through its chain-of-thought generation. GPT-4o shows high final-answer accuracy on straightforward problems but frequently omits intermediate steps and struggles with issues involving implicit constraints. In contrast, DeepSeek-V3 can reconstruct the multi-step problem-solving process more stably and performs well in problems centred on concept understanding. However, its effectiveness is significantly reduced when dealing with issues involving ambiguous language or implicit phrasing. Although each LLM demonstrated particular areas of strength, we also observed that they all have certain limitations in interpreting questions accurately, adapting reasoning flexibly, and maintaining logical completeness. These findings show that the reasonable design of prompt words and the addition of intermediate-step guidance are crucial to improving the robustness of the model.
Table 18: Effect of Prompt Design on LLMs Responses (Illustrative Cases)
Prompt Type Model Outcome Statement
Find the total cost GPT-4o incorrect Model failed to perform unit conversion and no intermediate steps shown.
Show all steps to calculate the total cost and unit conversions GPT-4o correct Model gave all clear steps and accurate calculations.
Solve the value of X DeepSeek incorrect Skipped variable substitution and used wrong formula.
State the formula and justify each step to find X DeepSeek correct Using the correct formula and giving the logical justification for each step.
Solve this problem Gemini Partially correct Answer included but reasoning was vaggue and generalised.
Show the each step for solving this problem and give the statement Gemini correct Provided well-structured reasoning with consistent logic.
Table 19: Qualitative Comparison of LLMs on the UNSW Dataset
Dimension GPT-4o DeepSeek Gemini
Final answer accuracy High Moderate Moderate
Step completeness Moderate Low High
Step validity High Moderate High
Intermediate calculation Moderate High High
Problem understanding Moderate Low High
Handing of complex logic High Low Moderate
The results indicate that GPT-4o has an acceptable level of performance when responding to the UNSW questions (Table 12). Specifically, it shows a high final-answer accuracy across simple and moderately difficult problems, achieving an average correctness rate of 82%. However, a detailed analysis of the SCoT framework reveals multiple problems. The reasoning process tends to "skip steps." Though the ultimate answer could be right, the logical coherence and consistency of the middle steps are lost. This is especially crucial for the case of applied problems because GPT-4o does not tend to include the implicit conditions set by the question, leading to low problem understanding and computational correctness scores, resulting in a low step completeness score of 57% (Table 14). In addition, it is less likely to have explicit equations and relations between the variables, hurting the interpretability of the solution approach. Notably, GPT-4o’s performance is sensitive to the formulation of the prompt. We observe notable improvements in the model’s logical coherence and reasoning completeness when prompts explicitly request step-by-step instructions or clarify assumptions, particularly on complex UNSW problems (see Table 19 and Table 18). Although the model can produce correct results, its reliability in educational contexts depends heavily on structured prompt scaffolding.
As shown in Tables 14 and 19, our qualitative analysis for the UNSW dataset shows that DeepSeek-V3 performs well on conceptual and definitional questions with a computational accuracy of 81% and step validity rate of 76%, which is superior to that of GPT-4o, and can reconstruct the mathematical solution chain. However, the question comprehension score drops to 64% when faced with problems with high linguistic ambiguity, making it challenging to identify implied structural changes or unit conversions in the stem. The model is more dependent on system prompts, and complex problems may lead to correct answers but incorrect reasoning, with the risk of misleading instructions. Gemini-2.0 has a balanced performance on all five dimensions, with a step completeness rate of 82% and an intermediate accuracy of 80 % (Table 14), and can generate a stable reasoning process. It handles complex problems well, maintaining semantic consistency and logical coherence, which leads to a drop in final accuracy to 66%.
Through an in-depth analysis of UNSW higher-level mathematical problems, we found that the design of prompts significantly impacts the reasoning performance of large language models (LLMs) in higher education scenarios. The questions in the UNSW dataset require open-mindness, an academic style of expression, and step-by-step solutions, which challenge the model’s linguistic comprehension and logical reasoning ability. We observe that there are differences in the way different models respond to the prompts. GPT-4o significantly improves the completeness of its reasoning chain when explicitly asked to answer in steps. Gemini maintains a high level of linguistic consistency and clarity of expression in response to structured questions. Finally, DeepDeek is more prone to logical jumps or erroneous results in response to a lack of derivation steps.
Recent studies have focused on evaluating the efficiency of LLMs in mathematical reasoning tasks and providing insights by constructing new benchmark tests. Mirzadeh et al. [138] and Deng et al.[139] revealed the limitations of these models in mathematical reasoning, and the potential advantages suggested by Chain-of-Thought (CoT) hint at potential advantages. The results of our research match these findings and apply the SCoT framework to model evaluation, focusing on multiple performance dimensions. In addition, this study explores the model’s adaptability in an educational setting by introducing UNSW higher-level course topics, providing insights for pedagogical applications.
This study has practical implications for teaching mathematics at colleges and universities such as UNSW, and also applies to the rest of Australian universities. In traditional lectures and tutorials, the focus of assessment grading is qualitative analysis of the rigour of students’ reasoning in breaking two the problem and providing steps of how the solution has been reached. On the contrary, with well-designed hints, LLMs can automatically generate more complete solutions and provide hints for different types of errors (e.g., missing steps, unclear definitions, and computational errors), which can help improve the efficiency of instructional feedback. For example, in a statistical inference or optimisation course, the Lecturer can set up a standard prompt template to guide the model in producing a structured solution path in JSON format for students’ self-checking and assessment. The SCoT framework used in this study can be extended to the classroom as a teaching assessment tool. For example, ’completeness of steps’, ’accuracy of intermediate calculations’, and ’problem comprehension’ can be used as scoring dimensions to assess students’ submitted assignments or model-generated process answers. This way, a ’human-machine co-assessment’ mechanism can be constructed to achieve a more transparent and detailed teaching feedback system.
In future studies, we suggest integrating LLM into the teaching platform to build an intelligent teaching and learning system with subject suitability and role division of labour. Educators (lecturers/tutors/teachers) can preset the prompting strategies and assessment criteria for different course modules(e.g., probabilistic inference, matrix decomposition, financial modelling), and the system will call on various models to complete the explanation, training and correction tasks. This nested design, based on ’course structure + model characteristics + prompting strategies’, is expected to enhance the application value of LLM in university mathematics courses and promote the two-way improvement of personalised learning and teaching efficiency.

### v2

@@ Abstract
Large Language Models (LLMs) have shown impressive performance on a range of educational tasks, but are still understudied for their potential to solve mathematical problems. In this study, we compare three prominent LLMs, including GPT-4o, DeepSeek-V3, and Gemini-2.0, on three mathematics datasets of varying complexities (GSM8K, MATH500, and University datasets). We take a five-dimensional approach based on the Structured Chain-of-Thought (SCoT) framework to assess final answer correctness, step completeness, step validity, intermediate calculation accuracy, and problem comprehension. The results show that GPT-4o is the most stable and consistent in performance across all the datasets, but particularly it performs outstandingly in high-level questions of the UNSW dataset. DeepSeek-V3 is competitively strong in well-structured domains such as optimisation, but suffers from fluctuations in accuracy in statistical inference tasks. Gemini-2.0 shows strong linguistic understanding and clarity in well-structured problems but performs poorly in multi-step reasoning and symbolic logic. Our error analysis reveals particular deficits in each model: GPT-4o is at times lacking in sufficient explanation or precision; DeepSeek-V3 leaves out intermediate steps; and Gemini-2.0 is less flexible in mathematical reasoning in higher dimensions.
@@ 3 Methodology
We first present the framework used to evaluate selected LLMs for mathematical problem-solving. Our approach combines multiple benchmark datasets along with University assignments and exams. We incorporate controlled prompting strategies, with both quantitative and qualitative evaluation.
@@ 3.1 Mathematical problems
To ensure representativeness and reliability of the evaluation results, this study selected three sets of mathematical problems from different levels and backgrounds: GSM8K, MATH 500, and the graduate course question bank from a selected University. These sets cover a wide range of problem types, from basic arithmetic and algebraic operations to advanced statistical reasoning and optimisation, and are characterised by diverse structures and distinct language styles.
@@
    University Graduate Question Bank: This dataset consists of exam and assignment questions covering probability inference, matrix decomposition, optimisation theory, and other courses. The language is rigorous, and the expression style is academic. Some questions have open structures, conceptual explanations, and non-standard expressions. The details about this dataset can be shown in Table LABEL:tab:data_unsw.
@@
University Problem 130
@@ 3.2 Automated LLM Evaluation Framework
    University Problem dataset (real academic assessment scenarios (Table LABEL:tab:data_unsw))
@@
    Stage 2 - Dataset pre-processing: We select many mathematical questions from the GSM8K, MATH500, and University datasets (Table 1), covering multiple levels from basic arithmetic to advanced reasoning. After pre-processing, we classify the questions according to knowledge points and difficulty levels to facilitate subsequent layered evaluation and error analysis. We need to balance question types to ensure that different kinds of problems are represented in the evaluation.
@@ 3.3 Evaluation Metrics and LLM Prompting Strategies
The LLM evaluation focuses on two aspects: algebraic expression computation and mathematical reasoning problems [119]. Since LLMs are generative models, ensuring that the generated answers are accurate becomes very tricky, especially for mathematical reasoning tasks [120]. In the case of computational problems, we prefer to provide standard reference answers so that we can easily compare the model-generated answers to the correct answers. This facilitates the assessment of the model’s performance in terms of numerical accuracy. We adopt the automatic proof verification method for the mathematical proof problem. Instead of relying on manual evaluations, we test the performance of the model by establishing the SCoT framework and an automatic scoring method inspired by recent advances in LLM-based reasoning verification [ren2023self]. It provides a more objective criterion for evaluating mathematical reasoning problems and improves the efficiency of the evaluation process [122]. We also provide a prompting mechanism to assess the ability to correct their mistakes. If a model provided an incorrect answer, it would be prompted again to solve the problem, and the corrected answer would be evaluated.
@@ 4.1 LLMs for solving GSM8K problems
We investigated more about the errors contributing to this mediocre result, including interruptions in the reasoning chain, omissions of steps, misunderstandings of quantitative relationships, and errors in unit conversion. In the case of GSM8K, all three models sometimes arrives at the correct answer. However, they cannot generate intermediate steps due to an unstable reasoning process or improper handling of details, which reflects their lack of coherence and robustness in multi-step tasks, as shown in Figure LABEL:fig:sample_question_answer.
@@
Table 4 shows that Gemini performs relatively best on the GSM8K dataset, with an accuracy of 53.9%. In addition, it demonstrates a more complete reasoning path compared to GPT-4o (50.0%) and DeepSeek-V3 (42.6%)  [^2] https://github.com/sydney-machine-learning/LLM-mathematics. To better understand this difference, we qualitatively (manually) review representative model responses, as summarised in Table 5. Gemini produced fewer instances of step skipping, unit confusion, and misinterpretation of the question intent. In contrast, we also found that DeepSeek’s reasoning chain is the most unstable, with frequent errors in unit conversion and logical expression, and is at the bottom of the three in terms of stability and accuracy.